OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = 11x + 30?

OpenStudy (lgbasallote):

first put x^2 to the right side y = x^2 + 11x + 30 now put y to zero because it is an x-intercept 0 = x^2 + 11x + 30 does that look familiar?

OpenStudy (anonymous):

yes i think i understand

OpenStudy (lgbasallote):

great! ^_^

OpenStudy (anonymous):

so then -b/2a for the vertex, right?

OpenStudy (lgbasallote):

uhmm idk those formulas sorry...i just remember the basics haha..maybe try asking again in the main question and someone who remembers it can answer :)